OpenStudy (anonymous):

Solve the given equation to approximate the angle x. Round to three decimal places: 75sin^2x - 34sinx / 3 = 0

OpenStudy (anonymous):

if it helps, my choices are: x=0.116, 0.311 x=0.118, 0.322 x=0.120, 0.340 x=0.122, 0.358 x=0.124, 0.369

OpenStudy (anonymous):

it is:

OpenStudy (anonymous):

I wasn't sure whether that was a division symbol or a plus symbol

OpenStudy (anonymous):

how do i go from sinx=0, 34/225 to the angle x?

OpenStudy (kropot72):

@dpalnc The question is already a quadratic \[75\sin ^{2}x-34\sin x+3=0\]

OpenStudy (anonymous):

i'm not sure if the symbol is a + or a /

OpenStudy (anonymous):

thanks @kropot72 , that "+" really looks like a divide...

OpenStudy (anonymous):

ok, so we have sinx=0 and sinx=34/225. how do i then find angle x?

OpenStudy (anonymous):

the equation I have is wrong... after looking at your attachment, the equation is as @kropot72 says...

OpenStudy (anonymous):

ok, so starting over... what do i do?

OpenStudy (anonymous):

nevermind, i think i figured it out

OpenStudy (kropot72):

Please come back if you need help.